OpenStudy (anonymous):

Simplify 3 - (4x - 5) + 6. A) 12x + 21 Eliminate B) 4x + 14 C) 4x + 4 D) -4x + 14

OpenStudy (jahnavi):

(D) is the write answer

OpenStudy (anonymous):

3-(4x-5)+6=3-4x+5+6=-4x+14=D

OpenStudy (anonymous):

Thank you guys so much!

OpenStudy (anonymous):

distribute the negative, add together the like terms